OpenStudy (anonymous):

Identify the factors of x^2 + 36y^2. Prime (x + 6y)(x - 6y) (x + 6y)(x + 6y) (x - 6y)(x - 6y)

OpenStudy (anonymous):

I think it's A, but im not sure :(

OpenStudy (anonymous):

I feel like it could be B to.. because can't you factor 36 by 6?

OpenStudy (anonymous):

Any ideas anyone?:L

OpenStudy (anonymous):

\[(x-6 i y) (x+6 i y)=x^2+36 y^2 \]

OpenStudy (anonymous):

So B? @robtobey

OpenStudy (jhannybean):

Your answer choices are all wrong. Along with the question itself.

OpenStudy (anonymous):

That's what my homework gave me tho? .-.

OpenStudy (anonymous):

I'm pretty sure its B because of the formula you have to use.. a^2 - b^2 = (a + b)(a- b)

OpenStudy (jhannybean):

\[x^2+36y^2 \implies (x)^2+(6y)^2\implies ?\]x and 36y\(^2\) are perfect squares. You just need to find their base.

OpenStudy (anonymous):

b. is \[x^2+12 x y+36 y^2 \]

OpenStudy (anonymous):

So if none of the answering are up there it's prime?

OpenStudy (anonymous):

In the sense that it cannot be factored to give real numbers, yes.

OpenStudy (anonymous):

Thank you! @robtobey :)

OpenStudy (anonymous):

Your welcome and thank you for the medal.